OpenStudy (cheddarskittles):

.Show work? 8a - 4=10a - 14

Mehek (mehek14):

\(10-8=2\\2a\\14+-4=10\\10=2a\) divide 10/2 \(5=a\)

OpenStudy (cheddarskittles):

3(a + 1) - 5= 5a -2

Mehek (mehek14):

distribute \(3a-2=5a-2\\0=2a\\0=a\)